# Variational Control for Guidance in Diffusion Models

### v1

@@ Abstract
Diffusion models exhibit excellent sample quality, but existing guidance methods often require additional model training or are limited to specific tasks. We revisit guidance in diffusion models from the perspective of variational inference and control, introducing Diffusion Trajectory Matching (DTM) that enables guiding pretrained diffusion trajectories to satisfy a terminal cost. DTM unifies a broad class of guidance methods and enables novel instantiations. We introduce a new method within this framework that achieves state-of-the-art results on several linear and (blind) non-linear inverse problems without requiring additional model training or modifications. For instance, in ImageNet non-linear deblurring, our model achieves an FID score of 34.31, significantly improving over the best pretrained-method baseline (FID 78.07). We will make the code available in a future update.
@@ 1 Introduction
Figure 1: Our method guides diffusion sampling to fulfill external constraints. To this end, we optimize the local direction ${\mathbf{u}}_{t}^{*}$ via external constraints while respecting the original trajectory, see Eq. 11 (left, center). This recovers more accurate reconstructions accross tasks compared to classical guidance methods: Nonlinear deblurring (Right). Our method accurately captures most details, while competing methods introduce artifacts in the generated reconstructions.
@@
Consequently, guidance in diffusion models has received significant interest. However, the dominant approaches to classifier guidance (Dhariwal & Nichol, 2021) and classifier-free guidance (Ho & Salimans, 2021) require training additional models or retraining diffusion models for each conditioning task at hand, or are based on simplistic assumptions detrimental to sample quality (Kawar et al., 2022; Chung et al., 2022a; Song et al., 2022b; Pandey et al., 2024b).
In this work, we generalize classifier guidance to a variational control problem (Kappen, 2008). Inspired by ideas such as Control as Inference (Kappen et al., 2012; Levine, 2018), we model guided diffusion dynamics as a Markov chain with the control signals defined as variational parameters. Variational inference is applied to optimize the control signals, ensuring that the process satisfies the desired terminal conditions while keeping the generated samples close to the unconditional sample manifold, see Fig. 1a. We denote this framework as Diffusion Trajectory Matching (DTM).
Recent work on steering diffusion models has already incorporated ideas from optimal control (Huang et al., 2024; Rout et al., 2024). However, these works focus on a restricted class of control problems. This obscures the available design choices revealed through our novel framework. Indeed, we find that DTM generalizes and explicitly contains a large class of prior work on guidance. We demonstrate the utility of this generalization by introducing a new sampling algorithm that seamlessly integrates with state-of-the-art diffusion model samplers like DDIM (Song et al., 2022a) and adapts well to diverse downstream tasks.
@@
  • NDTM outperforms previous state-of-the-art baselines for solving challenging noisy linear and (blind) non-linear inverse problems such as superresolution and inpainting with diffusion models (Daras et al., 2024) on FFHQ-256 (Karras et al., 2021) and ImageNet-256 (Deng et al., 2009).
@@ 3.1 Variational Control for Diffusion Guidance
Choice of Terminal Cost. The terminal cost in Eq. 8 encodes desirable constraints on the final guided state. For instance, $\Phi({\mathbf{x}}_{0})\propto-\log p({\bm{y}}|{\mathbf{x}}_{0})$ could be the log-likelihood of a probabilistic classifier for class-conditional generation or the degradation process for solving inverse problems. For instance, Huang et al. (2024) adapt guidance for a non-differentiable terminal cost using a path integral control (Kappen, 2005) formulation. While an interesting direction for further work, we only assume that the terminal cost is differentiable for now.
@@
${\mathcal{C}}_{\text{tr}}$ $=\sum_{t}\mathbb{E}_{{\mathbf{x}}_{t}}[D_{\mathrm{KL}}\!\left(p({\mathbf{x}}_{t-1}|{\mathbf{x}}_{t},{\mathbf{u}}_{t})~{}\|~{}q({\mathbf{x}}_{t-1}|{\mathbf{x}}_{t})\right).$ (9)
Note that other divergence measures can be useful depending on the specific form of the diffusion posterior (Nachmani et al., 2021; Zhou et al., 2023; Pandey et al., 2024a; Holderrieth et al., 2024).
@@
Continuous-Time Variants. To apply DTM to continuous-time diffusion and flow matching, we adapt the transient costs ${\mathcal{C}}_{\text{tr}}$. We call the following Continuous-Time Diffusion Trajectory Matching (CT-DTM), derived for continuous-time diffusion (Song et al., 2020) in Section A.2:
@@ 3.2 Non-linear Diffusion Trajectory Matching (NDTM)
For the guided diffusion posterior, we choose the following parameterization such that the control can affect the diffusion dynamics non-linearly.
@@ 3.3 Specific Instantiations
Here, we present a simplified form of the NDTM objective in the context of DDIM (Song et al., 2022a).
@@
For the diffusion posterior parameterization in DDIM (Song et al., 2022a), the NDTM objective in Eq. 16 has the following tractable upper bound (see proof in Appendix A.1),
@@
The coefficients $\alpha_{t}$ and $\sigma_{t}$ are specific to DDIM (see Section A.1 for more details). Intuitively, the simplified NDTM loss in Eq. 17 measures the deviation between the guided and unguided dynamics, penalizing the magnitude of the control signal ${\mathbf{u}}_{t}$ (first term) and deviations in the noise predictions (second term). On the contrary, the terminal loss ensures that the expected final guided state satisfies the external constraints. Therefore, the first two terms in Eq. 17 act as regularizers on the control signal ${\bm{u}}_{t}$.
In Section A.2, we derive this simplification also for continuous-time diffusion models.
@@ 3.4 Connection to Existing Guidance Mechanisms
Notably, if the terminal cost takes the form of a classifier likelihood $\Phi({\mathbf{x}}_{0})\propto-\log p({\mathbf{y}}|{\mathbf{x}}_{0})$, it can be shown (Huang et al., 2024) that the optimal control simplifies to classifier guidance (Dhariwal & Nichol, 2021): ${\mathbf{u}}_{t}^{*}=g(t)w_{T}\nabla_{{\mathbf{x}}_{t}}p({\mathbf{y}}|{\mathbf{x}}_{t})$.
This puts a large class of methods approximating the expectation over the posterior $p({\mathbf{x}}_{0}|{\mathbf{x}}_{t})$ (Chung et al., 2022a; Song et al., 2022b; Pandey et al., 2024b; Huang et al., 2024) into perspective: In terms of our DTM framework, they perform optimal control with a linear control mechanism. Empirically, we will see in Section 4 that our generalization to non-linear control provides significant performance improvements.
@@ 4 Experiments
While our method serves as a general framework for guidance in diffusion models, here, we focus on solving inverse problems (IP). Through both quantitative and qualitative results, we demonstrate that our approach outperforms recent state-of-the-art baselines across various (blind) non-linear and linear inverse problems. Lastly, we emphasize key design parameters of our proposed method as ablations. We present full implementation details in Appendix B.
@@
Problem Setup Given a corruption model ${\mathcal{A}}$ and a noisy measurement $\mathbf{y}\in R^{d}$ the goal is to recover the unknown sample ${\mathbf{x}}_{0}\sim p_{\text{data}}$, from the degradation $\mathbf{y}={\mathcal{A}}({\bm{x}}_{0})+\sigma_{y}{\mathbf{z}},\quad{\mathbf{z}}\sim{\mathcal{N}}\left(\mathbf{0},{\bm{I}}_{d}\right)$. For linear inverse problems, $\mathbf{y}={\mathcal{A}}{\bm{x}}_{0}$ where ${\mathcal{A}}$ can be any matrix. In the case where only the functional form of the degradation operator ${\mathcal{A}}$ is known but its parameters are not, the problem is known as Blind IP.
@@
Models and Datasets: We conduct experiments on the FFHQ $(256\times 256)$ (Karras, 2019) and ImageNet $(256\times 256)$ (Deng et al., 2009) datasets, using a held-out validation set of $1,000$ samples from each. For FFHQ, we use the pre-trained model provided by Chung et al. (2022a), and for ImageNet, we use the unconditional pre-trained checkpoint from OpenAI (Dhariwal & Nichol, 2021).
@@
Tasks and Metrics: We consider two linear and two non-linear inverse problems with noise level $\sigma_{y}=0.01$. For linear inverse problems, we use inpainting and super-resolution. For non-linear problems, we perform non-linear deblurring and blind image deblurring (BID) (see Appendix B).
@@
To quantitatively evaluate our results, we report metrics optimized for perceptual quality, including Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al., 2018b), Fréchet Inception Distance (FID) (Heusel et al., 2017), and Kernel Inception Distance (KID) (Bińkowski et al., 2018). For completeness, recovery metrics like the Peak Signal-to-Noise Ratio (PSNR) are provided in Appendix C. With the exception of BID (for which we use 100 images), we evaluate all other inverse problems on 1k images.
Baselines For linear tasks, we compare our proposed method against several state-of-the-art approaches, namely DPS (Chung et al., 2022a), RED-diff (Mardani et al., 2023b), C-$\Pi$GDM (Pandey et al., 2024b), and DDRM (Kawar et al., 2022). Since the latter two are restricted to linear degradation settings, for non-linear deblurring, we only compare our method with DPS and RED-diff. We also consider a recently proposed method, DMPlug (Wang et al., 2024), as a potential baseline. However, we find that DMPlug is too computationally expensive to evaluate on 1k samples (see Appendix C for details on runtimes). Therefore, we only use this baseline for the BID task evaluated on 100 images from the FFHQ dataset since, to the best of our knowledge, it is currently state-of-the-art in this task.
Sampling Setup. We tune all samplers for the best perceptual quality. We highlight sampler parameters that are common across all tasks and discuss task-specific sampler configs in the Appendix B.
  • DPS: We adapt the tuned hyperparameters for DPS from (Mardani et al., 2023b). More specifically, we set DDIM $\eta=0.5$, number of sampling steps to 1000.
  • RED-diff: We set $\sigma_{0}=0$ with a linear weighting schedule and $lr=0.5$ and perform 50 diffusion steps. We tune $\lambda$ depending on the task.
  • DDRM: We fix $\eta=0.85$, $\eta_{b}=1.0$, and the number of diffusion steps to 20 across all linear IPs.
  • C-$\Pi$GDM: We set the number of diffusion steps to 20 since C-$\Pi$GDM (Pandey et al., 2024b) saturates in performance within this budget. Furthermore, since Pandey et al. (2024b) report performance on noiseless inverse problems, we re-tune C-$\Pi$GDM for best perceptual quality (with $\sigma_{y}=0.01$) on all tasks and provide detailed hyperparameters in the Appendix.
  • DMPlug: Following (Wang et al., 2024), we fix diffusion steps to $3$, number of optimization steps to $10$k, and the (kernel, method) learning rates to (0.01, 0.1).
  • (Ours) NDTM: We fix the learning rate for the control, ${\mathbf{u}}_{t}$, to 0.01 across all tasks and apply a linear decay schedule over optimization steps. We tune the guidance weight $\gamma$, loss weighting, and DDIM $\eta$ for best performance across different tasks. We choose the number of optimization steps $N$ based on a tradeoff between sample quality and compute.
4.1 Non-linear Inverse Problems
Non-Linear Deblurring. As a first non-linear IP, we consider non-linear debluring with the same setup as in Chung et al. (2022a). Figure 2 (Left) illustrates the comparison between competing baselines and our proposed method, NDTM, on this task. Qualitatively, similar to Mardani et al. (2023a), we find that DPS is very sensitive to guidance step size and is usually unstable on this task. Moreover, while RED-diff does not have stability issues, we find that it is biased towards generating blurry samples. This is not surprising given their unimodal approximation to the data posterior $p({\mathbf{x}}_{0}|{\mathbf{y}})$. On the contrary, NDTM generates high-fidelity reconstructions with a stable sampling process. Similarly, our quantitative results in Table 1 validate our qualitative findings as our method outperforms competing baselines on perceptual quality for both datasets by a significant margin.

### v2

@@ Abstract
Diffusion models exhibit excellent sample quality, but existing guidance methods often require additional model training or are limited to specific tasks. We revisit guidance in diffusion models from the perspective of variational inference and control, introducing Diffusion Trajectory Matching (DTM) that enables guiding pretrained diffusion trajectories to satisfy a terminal cost. DTM unifies a broad class of guidance methods and enables novel instantiations. We introduce a new method within this framework that achieves state-of-the-art results on several linear, non-linear, and blind inverse problems without requiring additional model training or specificity to pixel or latent space diffusion models. Our code will be available at https://github.com/czi-ai/oc-guidance.
@@ 1 Introduction
Consequently, guidance in diffusion models has received significant interest. However, the dominant approaches to classifier guidance (Dhariwal & Nichol, 2021) and classifier-free guidance (Ho & Salimans, 2021) require training additional models or retraining diffusion models for each conditioning task at hand, or are based on simplistic assumptions detrimental to sample quality (Kawar et al., 2022; Chung et al., 2022a; Song et al., 2023; Pandey et al., 2024b).
A trained diffusion model can be viewed as a steerable stochastic system that follows learned dynamics. By modifying its inputs, we can guide it in the spirit of stochastic optimal control — aiming to reach desired terminal states (e.g., matching a guidance signal) while staying close to its native trajectory distribution. This intuition motivates our formulation of classifier guidance as a variational control problem (Kappen, 2008). Inspired by Control as Inference frameworks (Kappen et al., 2012; Levine, 2018), we model guided diffusion as a controlled Markov process, where control signals are treated as variational parameters. We then apply variational inference to optimize these controls, ensuring that generated samples satisfy terminal constraints without straying far from the unconditional sample manifold (see Fig. 1a). We refer to this framework as Diffusion Trajectory Matching (DTM).
Recent work on steering diffusion models has already incorporated ideas from optimal control (Huang et al., 2024; Rout et al., 2025). However, these works focus on a restricted class of control problems. This obscures the available design choices revealed through our novel framework. Indeed, we find that DTM generalizes and explicitly contains a large class of prior work on guidance. We demonstrate the utility of this generalization by introducing a new sampling algorithm that seamlessly integrates with state-of-the-art diffusion model samplers like DDIM (Song et al., 2021a) and adapts well to diverse downstream tasks.
@@
  • We show that NDTM outperforms previous state-of-the-art baselines for solving challenging and diverse inverse problems with pretrained pixel-space (see Fig. 1) diffusion models. Furthermore, NDTM can be extended trivially to latent-space diffusion models like Stable Diffusion (Rombach et al., 2022) for tasks like style guidance (see Fig. 3).
@@ 2 Background
Figure 1: Our method guides diffusion sampling to fulfill external constraints. To this end, we optimize the local direction ${\mathbf{u}}_{t}^{*}$ via external constraints while respecting the original trajectory, see Eq. 11 (left, center). This recovers more accurate reconstructions across tasks compared to classical guidance methods: Nonlinear deblurring (Right). Our method accurately captures most details, while competing methods introduce artifacts in the generated reconstructions.
@@ 3.1 Variational Control for Diffusion Guidance
Choice of Terminal Cost. The terminal cost in Eq. 8 encodes desirable constraints on the final guided state. For instance, $\Phi({\mathbf{x}}_{0})\propto-\log p({\bm{y}}|{\mathbf{x}}_{0})$ could be the log-likelihood of a probabilistic classifier for class-conditional generation or the degradation process for solving inverse problems (see Section 5 for exact form of terminal costs for different tasks). For instance, Huang et al. (2024) adapts guidance for a non-differentiable terminal cost using a path integral control (Kappen, 2005) formulation. While an interesting direction for further work, we only assume that the terminal cost is differentiable for now.
@@
${\mathcal{C}}_{\text{tr}}$ $=\sum_{t}\mathbb{E}_{{\mathbf{x}}_{t}}\big{[}D_{\mathrm{KL}}\!\left(p({\mathbf{x}}_{t-1}|{\mathbf{x}}_{t},{\mathbf{u}}_{t})~{}\|~{}q({\mathbf{x}}_{t-1}|{\mathbf{x}}_{t})\right)\big{]}.$ (9)
Note that other divergence measures can be useful depending on the specific form of the diffusion posterior (Nachmani et al., 2021; Zhou et al., 2023; Pandey et al., 2025; Holderrieth et al., 2025).
@@
Continuous-Time Variants. To apply DTM to continuous-time diffusion and flow matching, we adapt the transient costs ${\mathcal{C}}_{\text{tr}}$. We call the following Continuous-Time Diffusion Trajectory Matching, derived for continuous-time diffusion (Song et al., 2020) in Section A.2:
@@ 3.2 Non-linear Diffusion Trajectory Matching (NDTM)
In analogy to how the unguided diffusion denoising process $q$ is parameterized, we define our controlled process $p$ as
@@ 3.3 Specific Instantiations
Here, we present a simplified form of the NDTM objective in the context of DDIM (Song et al., 2021a).
@@
For the diffusion posterior parameterization in DDIM (Song et al., 2021a), the NDTM objective in Eq. 16 has the following tractable upper bound (see proof in Appendix A.1),
@@
The coefficients $\alpha_{t}$ and $\sigma_{t}$ are specific to DDIM (see Section A.1 for more details). Intuitively, the simplified NDTM loss in Eq. 17 measures the deviation between the guided and unguided dynamics, penalizing the magnitude of the control signal ${\mathbf{u}}_{t}$ (first term) and deviations in the noise predictions (second term). On the contrary, the terminal loss ensures that the expected final guided state satisfies the external constraints. Therefore, the first two terms in Eq. 17 act as regularizers on the control signal ${\bm{u}}_{t}$. In Section A.2, we derive this simplification also for continuous-time diffusion models. Lastly, it is worth noting that the control loss in Eq. 17 can be generalized as,
${\mathcal{C}}({\mathbf{u}}_{t})=w_{c}\big{\|}{\mathbf{u}}_{t}\big{\|}_{2}^{2}+w_{s}\big{\|}\epsilon_{\theta}(\bar{{\mathbf{x}}}_{t},t)-\epsilon_{\theta}({\mathbf{x}}_{t},t)\big{\|}_{2}^{2}+w_{T}\Phi(\hat{{\mathbf{x}}}_{0}^{t})$ (18)
In this work, unless specified otherwise, we set $w_{s}=\tau_{t}^{2}$, $w_{c}=\kappa_{t}^{2}$. However, exploring alternative weighting schedules for different loss terms can be an interesting direction for further work.
@@ 3.4 Connection to Existing Guidance Mechanisms
Notably, if the terminal cost takes the form of a classifier likelihood $\Phi({\mathbf{x}}_{0})\propto-\log p({\mathbf{y}}|{\mathbf{x}}_{0})$, it can be shown (Huang et al., 2024) that the optimal control simplifies to classifier guidance (Dhariwal & Nichol, 2021): ${\mathbf{u}}_{t}^{*}=g(t)w_{T}\nabla_{{\mathbf{x}}_{t}}p({\mathbf{y}}|{\mathbf{x}}_{t})$. This puts a large class of methods approximating the expectation over the posterior $p({\mathbf{x}}_{0}|{\mathbf{x}}_{t})$ (Chung et al., 2022a; Song et al., 2023; Pandey et al., 2024b; Huang et al., 2024) into perspective: In terms of our DTM framework, they perform optimal control with a linear control mechanism (i.e. control added linearly to the score function). Empirically, we will see in Section 5 that our generalization to non-linear control provides significant performance improvements.
4 Related Work
Conditional Diffusion Models. In general, the conditional score $\nabla_{{\mathbf{x}}_{t}}\log p({\mathbf{x}}_{t}|{\bm{y}})$ needed for guided sampling can be learned during training (Saharia et al., 2022; Podell et al., 2024; Rombach et al., 2022) or approximated during inference (see Daras et al. (2024) for a detailed review).
Here, we focus on training-free guidance during inference. In this context, there has been some recent progress in approximating the noisy likelihood score (see Eq. 5) by approximating the diffusion posterior $p({\mathbf{x}}_{0}|{\mathbf{x}}_{t})$. For instance, DPS (Chung et al., 2022a) approximates the diffusion posterior by a Dirac distribution centered on Tweedie’s estimate (Efron, 2011). This has the advantage that the guidance can be adapted to linear and non-linear tasks alike. However, due to a crude approximation, DPS converges very slowly and, in our observation, could be unstable for certain tasks (see Table 1). Consequently, some recent work (Yu et al., 2023; Bansal et al., 2024) adds a correction term on top of the DPS update rule to better satisfy the constraints. MPGD (He et al., 2024) attempts to alleviate some of these issues by leveraging the manifold hypothesis. In contrast, our proposed method instead directly estimates the guided posterior at each sampling step, thus sidestepping the limitations of DPS in the first place.
It is worth noting that our method resembles DCPS (Janati et al., 2024), which adopts a similar approach of learning a series of potentials to sample from the final posterior. However, DCPS involves an additional overhead of Langevin Monte Carlo sampling in addition to the posterior optimization step. Moreover, the control perspective adopted in this work helps contextualize prior work in guidance within our framework (see Section 3.4), which is lacking in DCPS. Some recent methods like TFG (Ye et al., 2024) also propose general frameworks for guidance in diffusion models. However, our proposed method does not fit within their framework.
More recent work (Kawar et al., 2022; Wang et al., 2023; Song et al., 2023; Pandey et al., 2024b; Pokle et al., 2024; Boys et al., 2024) relies on expressive approximations of the diffusion posterior. While this can result in accurate guidance and faster sampling, a large proportion of these methods are limited to linear inverse problems, which further limits their application to pixel space diffusion models. In contrast, our method can be adapted to generic inverse problems and is thus agnostic to the diffusion model architecture. Lastly, another line of work in inverse problems approximates the data posterior $p({\mathbf{x}}_{0}|{\mathbf{y}})$ using variational inference (Blei et al., 2017; Zhang et al., 2018a). For instance, RED-diff (Mardani et al., 2024) proposes to learn an unimodal approximation to the data posterior by leveraging a diffusion prior. However, this can be too restrictive in practice and comes at the expense of blurry samples. We refer interested readers to Daras et al. (2024) for a more detailed review of training-free methods for solving inverse problems in diffusion models.
Optimal Control for Diffusion Models. There has been some recent interest in exploring connections between stochastic optimal control and diffusion models (Berner et al., 2024). Chen et al. (2024) leverages ideas from control theory for designing efficient diffusion models with straight-line trajectories in augmented spaces (Pandey & Mandt, 2023). Since our guided sampler can be used with any pretrained diffusion models, our approach is complementary to this line of work. More recently, SCG (Huang et al., 2024) leverages ideas from path integral control to design guidance schemes with non-differentiable constraints. In contrast, we only focus on differentiable terminal costs, and extending our framework to non-differentiable costs could be an important direction for future work. Lastly, Rout et al. (2025) propose RB-Modulation, a method based on control theory for personalization using diffusion models. Interestingly, while RB-Modulation is primarily inspired by a class of tractable problems in control theory, it is a special case of our framework in the limit of $w_{s}=0,w_{c}=0$ and $\gamma=1$. Therefore, our proposed framework is more flexible.
@@ 5 Experiments
While our method serves as a general framework for guidance in diffusion models, here, we focus on solving inverse problems and style-guided generation. Through both quantitative and qualitative results, we demonstrate that our approach outperforms recent state-of-the-art baselines across these tasks using pretrained diffusion models. Lastly, we emphasize key design parameters of our proposed method as ablations. We defer all implementation details to App. B.
@@
Problem Setup and Terminal Costs: For inverse problems, given a corruption model ${\mathcal{A}}$ and a noisy measurement $\mathbf{y}\in R^{d}$ the goal is to recover the unknown sample ${\mathbf{x}}_{0}\sim p_{\text{data}}$, from the degradation $\mathbf{y}={\mathcal{A}}({\bm{x}}_{0})+\sigma_{y}{\mathbf{z}},\quad{\mathbf{z}}\sim{\mathcal{N}}\left(\mathbf{0},{\bm{I}}_{d}\right)$. For linear inverse problems, $\mathbf{y}={\mathcal{A}}{\bm{x}}_{0}$. In the case where only the functional form of the degradation operator ${\mathcal{A}}$ is known but its parameters are not, the problem is known as Blind Inverse Problem. For inverse problems, we consider the following form of the terminal cost $\Phi(\hat{{\mathbf{x}}}_{0}^{t})$ in Eq. 17,
$\Phi(\hat{{\mathbf{x}}}_{0}^{t})=\|{\bm{y}}-{\mathcal{A}}(\hat{{\mathbf{x}}}_{0}^{t})\|_{2}^{2}$ (21)
where $\hat{{\mathbf{x}}}_{0}^{t}$ is the Tweedies estimate at any given time t. We also consider the task of style guidance with Stable Diffusion, where the goal is to generate samples adhering to a specific prompt and a reference style image. More specifically, given a reference style image ${\bm{r}}$, a pretrained feature extractor (like CLIP) ${\mathcal{F}}$, a pretrained decoder ${\mathcal{D}}$ such that $\hat{{\mathbf{x}}}_{0}^{t}={\mathcal{D}}(\hat{{\mathbf{z}}}_{0}^{t})$, we define $\Phi(\hat{{\mathbf{x}}}_{0}^{t})$ as,
$\Phi(\hat{{\mathbf{x}}}_{0}^{t})=\|{\mathcal{G}}({\mathcal{F}}({\bm{r}}))-{\mathcal{G}}({\mathcal{F}}(\hat{{\mathbf{x}}}_{0}^{t}))\|_{F}^{2}$ (22)
where ${\mathcal{G}}$ denotes the Gram-matrix operation and .$\|.\|_{F}$ denotes the Frobenius norm. Note that prompt adherence is achieved via the pretrained Stable Diffusion model.
@@
Models and Datasets: For inverse problems, we conduct experiments on the FFHQ $(256\times 256)$ (Karras et al., 2019) and ImageNet $(256\times 256)$ (Deng et al., 2009) datasets, using a held-out validation set of $1,000$ samples from each. For FFHQ, we use the pre-trained model provided by Chung et al. (2022a), and for ImageNet, we use the unconditional pre-trained checkpoint from OpenAI (Dhariwal & Nichol, 2021). For style guidance, following MPGD (He et al., 2024), we randomly generate 1k (prompt, image) pairs using images from WikiArt (Saleh & Elgammal, 2015) and prompts from PartiPrompt (Yu et al., 2022). We use the pre-trained CLIP encoder and Stable Diffusion 1.4 models as the feature extractor ${\mathcal{F}}$ and diffusion model, respectively.
@@
Tasks and Metrics: For inverse problems, we consider random inpainting, super-resolution, and non-linear deblurring. Additionally, we consider blind image deblurring (BID) task where we additionally infer the deblurring kernel ${\bm{k}}$ along with the final reconstruction. We set the noise level $\sigma_{y}=0.01$ for all inverse problems.
@@
For quantitative evaluation on inverse problems, we report metrics optimized for perceptual quality, including Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al., 2018b), Fréchet Inception Distance (FID) (Heusel et al., 2017), and Kernel Inception Distance (KID) (Bińkowski et al., 2018). For completeness, recovery metrics like the Peak Signal-to-Noise Ratio (PSNR) are provided in Appendix C. With the exception of BID (for which we use 100 images), we evaluate all other inverse problems on 1k images. For style guidance, following prior work Yu et al. (2023); He et al. (2024), we report the CLIP score (which measures prompt adherence) and the Style Score (which measures style adherence) on 1k text and image pairs.
5.1 Inverse Problems
We first evaluate the proposed NDTM sampler against competing baselines for non-linear, blind, and linear inverse problems. We provide all hyperparameter details for our method and competing baselines in Appendix B.